OpenStudy (kevin):

If the operation (a mod b) is the remainder of division operation by b, calculate (7^7.777.777 mod 100) + (5^5.555.555 mod 10)? The answer is 12. Pls tell me step by step how to get its number.

OpenStudy (kevin):

@mathmate @mathmale @skullpatrol @RadEn @princeharryyy

OpenStudy (unavailabilityy):

Um, what grade is this? xD

OpenStudy (kevin):

arithmetic for computer science XD

OpenStudy (unavailabilityy):

Uhh

OpenStudy (kevin):

yup

OpenStudy (princeharryyy):

what are dots in 7777777

OpenStudy (mathmate):

It just takes a little patience! Make a table, I'll start witht he first 3 entries, and will let you finish it until you find a pattern. n 7^n mod(7^n,100) 1 7 7 2 49 49 3 343 43 4 ... If the pattern has a period of k, then 7^7777777\(\equiv\)7^(7777777 mod k), the latter of which can be calculated using a calculator. You can do the same with 5^5555555 mod 10. I have verified the answer, which is indeed 12.

OpenStudy (kevin):

I'm sorry. Dots is not decimal symbol.

OpenStudy (princeharryyy):

okay

OpenStudy (princeharryyy):

it can be done manually.

OpenStudy (princeharryyy):

amd yes the answer is right.

OpenStudy (princeharryyy):

it is 12

OpenStudy (kevin):

the pattern will recur every 4x. Then?

OpenStudy (princeharryyy):

5 to power anything will leave 5 in the end... see, 5^1 =5 5^2 = 25 5^3 = 125 and so on. and because we have to find the mod we are concerned with last digit of 5^ whatever. So the remainder of second part would be 5. Right?

OpenStudy (kevin):

I know how to find the remainder. So what?

OpenStudy (princeharryyy):

you are looking for remainder in this question or are you not?

OpenStudy (kevin):

I mean after find the remainder. What we are going to do

OpenStudy (mathmate):

So you found k=4, Next step is to use 7^7777777 mod 100 ≡7^(7777777 mod k) mod 100

OpenStudy (princeharryyy):

we are going to find the remainder for the first one. LOL. We have already done the second part.

OpenStudy (princeharryyy):

when u divide the power 7777777 you would be left with only 1 power every thing else qould be contained in the division by 4.

OpenStudy (kevin):

Owhh... I get it now

OpenStudy (mathmate):

@Kevin You need to find the remainder of the each of 7^7777777 mod 100 and 5^5555555 mod 10. After that, the sum should equal 12.

OpenStudy (kevin):

It would be 7^1 mod 100 + 5^1 mod 100 ?

OpenStudy (princeharryyy):

7^7777777 = 7^1

OpenStudy (princeharryyy):

7^1 mod 100 + 5^1 mod 10 = 7 +5 =12

OpenStudy (kevin):

Sorry, that's what I mean 7^1 mod 100 + 5^1 mod 10

OpenStudy (kevin):

Thx guys

OpenStudy (princeharryyy):

yes it's right.@Kevin

OpenStudy (kevin):

I don't know whom I give the medal since both of you help me. lol